You currently owe $2,500. If the bank charges 16.44% compounded semiannually and you want to pay off your debt in 3 years, how much must you pay every six months?
Answer #1

This question requires application of PV of an annuity, according to which

PV =

For this question,

r = 16.44%/2 = 8.22% (semi-annually), n = 3 * 2 = 6 semi-annual periods

2500 = P * 4.5922

P = $544.40